OpenStudy (anonymous):

Integrate from 1-4: 3( ln x /x^2)

OpenStudy (anonymous):

\[3\int\limits_{1}^{4}(lnx/x^2)dx\]

OpenStudy (turingtest):

integration by parts

OpenStudy (anonymous):

integration by parts?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

x ln x - x

OpenStudy (turingtest):

that's not the right answer I don't think...

OpenStudy (anonymous):

u = lnx du = 1/x dx dv = x^-2 v=-x^-1

OpenStudy (anonymous):

no,

OpenStudy (anonymous):

maybe.

OpenStudy (turingtest):

that's what I did...

OpenStudy (anonymous):

should work then, try that krystal

OpenStudy (anonymous):

ok. thanks